OpenStudy (anonymous):

The variables x and y vary inversely. Which equation relates x and y when x = -1 and y = -9?

OpenStudy (anonymous):

K=1 OVER 9

OpenStudy (anonymous):

then what do I do?..

OpenStudy (anonymous):

We're supposed to try at explain @JULIANCHICO .. but a variables that vary inversely have the formula \[y = \frac{ k }{ x^2 }\] With this being said, we fill in the numbers next to find what K is equal to. \[-9 = \frac{ k }{ 1 }\] multiply both sides by 1 to get rid of the fraction. K = -9. So the answer you want is: \[y = \frac{ -9 }{ x^2 }\]

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

but none of the answers have x^2 in it..

OpenStudy (anonymous):

what are your choices?

OpenStudy (anonymous):

x = -y/9, y=-9x, x=9y, y=-9/x

OpenStudy (anonymous):

ohh okay I found my mistake, I just assumed it was x^2 like what I have in algebra 2 at the moment. Hang on let me re-do it real quick.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Interesting..when I re-do it I keep getting y = 9/x

OpenStudy (anonymous):

hmm

OpenStudy (mertsj):